# Suppression of the type Ia supernova host galaxy step in the outer regions of galaxies

### v1

@@ 4 Discussion
We have uncovered a relationship between the Hubble residuals of SNe Ia and the projected, normalized galactocentric distance of the SN from its host galaxy. After standardization for stretch and colour, SNe Ia in the inner regions of their host galaxies ($d_{\mathrm{DLR}}\leq 0.5$) have more negative Hubble residuals (i.e., appear brighter) than those at moderate ($0.5<d_{\mathrm{DLR}}\leq 1$) galactocentric distances (Fig. 4). More significantly, we have found that the magnitude of the step in Hubble residuals that occurs at stellar masses around $10^{10}\,\mathrm{M}_{\sun}$ is dependent upon the galactocentric distance: the step is large for SNe Ia with $d_{\mathrm{DLR}}\leq 1$ but negligible for SNe Ia with $d_{\mathrm{DLR}}>1$ (Fig. 5).
@@ 4.1 Dust
Previous studies have proposed an explanation of the host mass step in the form of a change in the ratio of total-to-selective extinction of the dust ($R_{V}$) in low and high-mass galaxies (Brout & Scolnic, 2021). A key motivation for the introduction of the Brout & Scolnic (2021) model is that it can naturally explain the observation that the size of the mass step increases for redder SNe Ia. This trend between mass step size and SN colour is present (before the BBC 4D bias correction) in DES-SN5YR data (Vincenzi et al., 2024), as well as the Pantheon+ (Brout et al., 2022) and Amalgame (Popovic et al., 2024b) compilations, although the trend is not significant in the low-$z$ ZTF sample (Ginolin et al., 2024b).
@@ 4.1.1 SN colour distributions
We perform two-sample statistical tests to determine whether the colour distributions are different. The SALT3 $c$ distributions for SNe in inner ($d_{\mathrm{DLR}}\leq 1$) and outer ($d_{\mathrm{DLR}}>1$) regions (Fig. 9) are consistent with being sampled from the same parent distribution: the two-sided Kolmogorov–Smirnov (KS) test statistic is 0.05 with a $p=0.29$, while a Mann–Whitney U test results in $p=0.12$. However, when splitting instead at $d_{\mathrm{DLR}}=1.5$, we find with high significance that the samples come from different populations (KS $p=5.6\times 10^{-4}$, Mann–Whitey $U$ $p=1.9\times 10^{-3}$).
To test whether this effect is caused by selection effects, we define a ‘colour test’ sample. We limit the redshift to $z<0.6$ to retain higher quality light-curves, while relaxing the SN colour selection of the cosmological sample and allowing the reddest SNe Ia up to $c<0.5$ to probe the full reddening tail. This gives 1268 SNe Ia. We repeat the tests, and the results are similar after having made this selection.
Interestingly, if instead we split the samples at $d_{\mathrm{DLR}}=1.5$, we find that the samples are significantly different, both in the nominal DES-SN5YR and the ‘colour test’ selections. This result indicates that SNe Ia at $d_{\mathrm{DLR}}>1.5$ are less affected by reddening, and is consistent with the findings of Ginolin et al. (2024b) [^2] Note that DLR is defined differently between ZTF and DES, with ZTF defining DLR=1 as the half-light radius, which is roughly DLR=0.68 in DES; thus the ZTF choice of $d_{\mathrm{DLR}}>1.5$ for dustless SNe Ia is consistent with our split of $d_{\mathrm{DLR}}=1$., but does not explain why $d_{\mathrm{DLR}}=1$ is the point at which there is the most significant difference in mass steps.
@@
The results are presented in Table 3, and the example of the model fitted to the ‘colour test’ sample is shown in Fig. 9. As above, when splitting the DES-SN5YR sample at $d_{\mathrm{DLR}}=1$ we find $\mu_{c}$, $\sigma_{c}$ and $\tau_{E}$ are consistent between SNe Ia in the inner and outer regions, inconsistent with the results from Ginolin et al. (2024b) where $\tau_{E}$ is reduced by $\sim 30$ per cent in the outer regions. We find consistent parameters even when limiting to the ‘colour test’ sample. Once more, if we split at $d_{\mathrm{DLR}}=1.5$ we find $\tau_{E}$ is significantly larger in outer regions, with the difference between inner and outer regions enhanced in the ‘colour test’ sample. No matter our selection, the DES data are not fit by as large a $\tau_{E}$ as the low-$z$ ZTF Ginolin et al. (2024b) sample, due to selection effects, choice of light curve fitter, and possible redshift evolution. Therefore direct comparisons between the colour distributions are not meaningful.
@@ 4.1.2 Hubble residuals as a function of SN Ia colour
Given the consistency of the SN Ia colour distributions between SNe Ia in inner and outer regions (Section 4.1.1), we now consider the slope of the dust extinction laws as a possible reason for the difference in step size.
The Hubble residuals are plotted against SN Ia colour for inner and outer regions in the left and right hand panels of Fig. 10 respectively. In inner regions, the mass step for all colours redder than $c=-0.1$ is clear, and increases with increasing colour: this is consistent with what was seen in the full DES-SN5YR sample (Kelsey et al., 2023; Vincenzi et al., 2024). The step in blue SNe Ia is not explained well by differing $R_{V}$ distributions and has been attributed to SNe Ia from differing ages of stellar population (Kelsey et al., 2023; Wiseman et al., 2022; Popovic et al., 2024a). The step in blue SNe Ia in inner regions, but not in outer regions, may be explained by a more diverse stellar (and SN progenitor) population in the inner regions of galaxies, while the outer regions host a homogeneous population of SNe Ia regardless of the host mass. We can invoke a similar description for dust: inner regions evolve such that their average dust laws are different for host galaxies of different mass, while outer regions retain the same dust properties throughout the evolution of the galaxy as a whole.
@@ 4.2 Populations of SN Ia progenitors
A difference in the effects of dust extinction between inner and outer regions does not seem to be the cause of the difference in host galaxy steps between those regions. A different explanation could be that the SN explosions themselves occur in multiple progenitors, with a different average standardized peak brightness. These populations are then present in differing proportions between the inner regions of low-mass/blue/young and high-mass/red/old galaxies, but the outer regions of all galaxies only host a single population.
These progenitors are often assumed to relate to the ‘delay time’ from an initial burst of star formation to the SN Ia explosion. For the SN Ia population as a whole, the distribution of these delay times – the delay time distribution (DTD) – is well described by a declining power law (Maoz & Gal-Yam, 2004; Totani et al., 2008; Maoz et al., 2010, 2014; Graur et al., 2015; Castrillo et al., 2021; Wiseman et al., 2021). When such a DTD is convolved with the stellar age distribution of stars across different galaxy types, a bimodal distribution for the age of SN progenitors at the time of explosion is observed (Mannucci et al., 2005; Scannapieco & Bildsten, 2005; Mannucci et al., 2006; Sullivan et al., 2006; Smith et al., 2012; Childress et al., 2014; Wiseman et al., 2021): so-called ‘prompt’ and ‘delayed ’ components.
@@
We fit the full dataset with this model by minimizing the negative log-likelihood, and show the data and best-fitting model in Fig 13. We find $\mu_{1,x_{1}}=0.22\pm 0.03$, $\sigma_{1,x_{1}}=0.67\pm 0.03$, $\mu_{2,x_{1}}=-0.86\pm 0.17$, $\sigma_{2,x_{1}}=0.80\pm 0.07$, and $a=0.37\pm 0.1$ which are broadly consistent with the results of Nicolas et al. (2021) and Ginolin et al. (2024a) except for $a$ which is significantly smaller than found in those works but consistent with the value for DES-SN5YR modelled in Wiseman et al. (2022). This small $a$ could be caused by our fixing of $y^{i}$ based on the galaxy colours: it is known that no galaxy observable perfectly traces the underlying properties, and even locally measured colours experience contamination whereby young SNe occur in populations measured to be old (Briday et al., 2022). If instead we fix the probability that any given SN is young to the average for our mean redshift according to the Rigault et al. (2020) drift model, $\delta(z=0.53)=0.67$, we find $\mu_{1,x_{1}}=0.06$, $\sigma_{1,x_{1}}=0.70$, $\mu_{2,x_{1}}=-1.65$, $\sigma_{2,x_{1}}=0.45$, and $a=0.58$. Note that these mean values are slightly more negative than the low-$z$ values from Nicolas et al. (2021) and Ginolin et al. (2024a) – this could be genuine redshift evolution or an effect of the use of SALT2 rather than SALT3 light curve fits in those works.
Next, we fit the inner and outer SNe separately, fixing the model parameters to those found using the aperture/annulus colour as a proxy for stellar population age, except for $a$ which we leave free. We find $a=0.32\pm 0.05$ in inner regions and $a=0.42\pm 0.04$ in outer regions, which means there is a larger fraction of high-stretch mode SNe in red outer annuli compared to red central regions.
@@ 5 Summary
Using the DES five-year sample of photometrically-classified SNe Ia, we have analysed the effects of projected and normalized host galaxy separation on the light-curve properties and inferred cosmological distance measurements. Using the directional light radius (DLR) method as our normalisation of the galactocentric distances ($d_{\mathrm{DLR}}$), our main findings are:
  • We confirm previous findings of Toy et al. (2023) that the innermost regions of galaxies ($d_{\mathrm{DLR}}$$\leq 1$) host faster-declining SNe than the outer regions.
  • We show, for the first time, that the difference in SN Ia post-standardization brightnesses between high and low-mass hosts reduces from $0.078\pm 0.011$ mag in the full sample to $0.036\pm 0.018$ mag for SNe Ia located in the outer regions of their host galaxies ($d_{\mathrm{DLR}}>1$), while increasing to $0.100\pm 0.014$ mag for SNe in the inner regions ($d_{\mathrm{DLR}}\leq 1$). The effect remains when splitting SNe Ia by their global galaxy $U-R$ colour, or by the $U-R$ colour in the inner aperture/outer annulus in which the SN occurred.
@@
  • There is no evidence that the outer regions of galaxies have dust laws that change as a function of stellar mass. Similarly, there is no evidence for an intrinsic luminosity difference between SNe Ia in the outer regions of low and high-mass galaxies.
@@
Calculating $d_{\mathrm{DLR}}$ is an algorithmically straight forward task, and selecting SNe Ia based on $d_{\mathrm{DLR}}$ reduces the need to account for host galaxy properties to standardize SN Ia brightnesses across different galaxies, without much computational effort. While we have not identified the underlying astrophysics that such projected galactocentric distances are tracing in the DES five-year SN Ia sample, we have shown that the standardized distance measurements from SNe Ia in the outer regions of galaxies have little dependence on their global host galaxy properties.
Restricting a cosmological analysis to SNe Ia in the outer regions of their host galaxies reduces the sample size to around a third. This will increase the statistical uncertainties. However, the reduction in the astrophysical systematic uncertainties (and complications) gained from using such a sample, coupled with the very large sample sizes expected in future experiments such as the Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), means that such a selection is likely to be beneficial.
@@ Acknowledgements
All authors have contributed to the drafting of this manuscript. MT devised the project and led the analysis. PW and MS provided scientific guidance and support with the production of the manuscript. MS contributed host galaxy fitting and wrote substantial sections of the manuscript, and PW ran 1D bias corrections and provided the colour and stretch analyses. DS and MV internally reviewed the work and provided extensive feedback. CF, CL, JL, LK, LG, RK, PS, and TD provided comments on the analysis and interpretation. All aforementioned authors as well as AM, BP, BS, DB, and MS contributed to the DES-SN5YR data and methods used in this paper. The remaining authors have made contributions to this paper that include, but are not limited to, the construction of DECam and other aspects of collecting the data; data processing and calibration; developing broadly used methods, codes, and simulations; running the pipelines and validation tests; and promoting the science analysis.

### v2

@@ 4 Discussion
We have uncovered a relationship between the Hubble residuals of SNe Ia and the projected, normalized galactocentric distance of the SN from its host galaxy. After standardization for stretch and colour, SNe Ia in the inner regions of their host galaxies ($d_{\mathrm{DLR}}\leq 0.5$) have more negative Hubble residuals (i.e., appear brighter) than those at moderate ($0.5<d_{\mathrm{DLR}}\leq 1$) galactocentric distances (Fig. 4). More significantly, we have found that the magnitude of the step in Hubble residuals that occurs at stellar masses around $10^{10}\,\mathrm{M}_{\sun}$ is dependent upon the galactocentric distance: the step is large for SNe Ia with $d_{\mathrm{DLR}}\leq 1$ but negligible for SNe Ia with $d_{\mathrm{DLR}}>1$ (Fig. 5). This result adds to the growing body of evidence that the size mass step can be reduced by a variety of methods and sample selections (Brout & Scolnic, 2021; Boone et al., 2021).
@@ 4.1 Dust
Previous studies have proposed an explanation of the host mass step in the form of a change in the ratio of total-to-selective extinction of the dust ($R_{V}$) in low and high-mass galaxies (Brout & Scolnic, 2021). A key motivation for the introduction of the Brout & Scolnic (2021) model is that it can naturally explain the observation that the size of the mass step increases for redder SNe Ia. This trend between mass step size and SN colour is present (before the BBC 4D bias correction) in DES-SN5YR data (Vincenzi et al., 2024), as well as the Pantheon+ (Brout et al., 2022) and Amalgame (Popovic et al., 2024b) compilations. However, Ginolin et al. (2024b) found no such significant trend between SN Ia Hubble residual and SN Ia colour in the low-$z$ ZTF sample.
@@ 4.1.1 SN colour distributions
We perform two-sample statistical tests to determine whether the colour distributions are different. The SALT3 $c$ distributions for SNe in inner ($d_{\mathrm{DLR}}\leq 1$) and outer ($d_{\mathrm{DLR}}>1$) regions are consistent with being sampled from the same parent distribution: the two-sided Kolmogorov–Smirnov (KS) test statistic is 0.05 with a $p=0.22$, while a Mann–Whitney U test results in $p=0.10$. However, when splitting instead at $d_{\mathrm{DLR}}=1.5$, we find mild evidence that the samples come from different populations (KS $p=0.03$, Mann–Whitey $U$ $p=0.008$).
To test whether this effect is caused by selection effects, we define a ‘colour test’ sample. We limit the redshift to $z<0.6$ to retain higher quality light-curves, while relaxing the SN colour selection of the cosmological sample and allowing the reddest SNe Ia up to $c<0.5$ to probe the full reddening tail. This sample is shown in Fig. 9. This gives 1268 SNe Ia. We repeat the tests: the two-sided KS test statistic is 0.05 with a $p=0.29$, while a Mann–Whitney U test results in $p=0.12$. However, when splitting instead at $d_{\mathrm{DLR}}=1.5$, we find with high significance that the samples come from different populations (KS $p=5.6\times 10^{-4}$, Mann–Whitey $U$ $p=1.9\times 10^{-3}$).
The result that the colour distributions appear different when split at $d_{\mathrm{DLR}}=1.5$, but not 1, both in the nominal DES-SN5YR and the ‘colour test’ selections, is interesting. This result indicates that SNe Ia at $d_{\mathrm{DLR}}>1.5$ are less affected by reddening, and is consistent with the findings of Ginolin et al. (2024b), but does not explain why a split of $d_{\mathrm{DLR}}=1$ leads to a more significant difference in mass steps.
@@
The results are presented in Table 3, and the example of the model fitted to the ‘colour test’ sample is shown in Fig. 9. We note that on visual inspection the modes of the model differ from the data, and differ between samples. The plotted model requires the addition of an uncertainty “smear" which we approximate by taking the median uncertainty of the data, but inevitably this leads to a slightly inaccurate representation of how the model fits the data. As above, when splitting the DES-SN5YR sample at $d_{\mathrm{DLR}}=1$ we find $\mu_{c}$, $\sigma_{c}$ and $\tau_{E}$ are consistent between SNe Ia in the inner and outer regions, which is different to what was found in Ginolin et al. (2024b) where $\tau_{E}$ is reduced by $\sim 30$ per cent in the outer regions. When limiting to the ‘colour test’ sample, we again find that the model parameters are consistent between inner and outer regions when split at $d_{\mathrm{DLR}}=1$, although for all $d_{\mathrm{DLR}}$ we find a significant shift to a redder $\mu_{c}$ and larger $\tau_{E}$ when compared to the full sample. However, if we split at $d_{\mathrm{DLR}}=1.5$ we find $\tau_{E}$ is significantly larger in outer regions, with the difference between inner and outer regions enhanced in the ‘colour test’ sample. No matter our selection, the DES data are not fit by as large a $\tau_{E}$ as the low-$z$ ZTF Ginolin et al. (2024b) sample, due to selection effects, choice of light curve fitter, and possible redshift evolution.
@@ 4.1.2 Hubble residuals as a function of SN Ia colour
We next consider the slope of the dust extinction laws as a possible reason for the difference in step size. The Hubble residuals are plotted against SN Ia colour for inner and outer regions in the left and right hand panels of Fig. 10 respectively. Note that the apparent increase in blue SNe to a positive Hubble residual is a well-known selection bias (see discussion in Scolnic & Kessler, 2016) resulting from our use of a BBC 1D bias correction.
In inner regions, the mass step for all colours redder than $c=-0.1$ is clear, and increases with increasing colour: this is consistent with what was seen in the full DES-SN5YR sample (Kelsey et al., 2023; Vincenzi et al., 2024). The step in blue SNe Ia is not explained well by differing $R_{V}$ distributions and has been attributed to SNe Ia from differing ages of stellar population (Kelsey et al., 2023; Wiseman et al., 2022; Popovic et al., 2024a). The step in blue SNe Ia in inner regions, but not in outer regions, may be explained by a more diverse stellar (and SN progenitor) population in the inner regions of galaxies, while the outer regions host a homogeneous population of SNe Ia regardless of the host mass. We can invoke a similar description for dust: inner regions evolve such that their average dust laws are different for host galaxies of different mass, while outer regions retain the same dust properties throughout the evolution of the galaxy as a whole.
@@ 4.2 Populations of SN Ia progenitors
A difference in the effects of dust extinction between inner and outer regions does not seem to be the cause of the difference in host galaxy steps between those regions. A different explanation could be that the SN explosions themselves occur in multiple types of progenitors, with a different average standardized peak brightness. These varied populations of progenitor systems could then be present in differing proportions between the inner regions of low-mass/blue/young and high-mass/red/old galaxies, but the outer regions of all galaxies only host a single population.
These progenitors types are often assumed to relate to the ‘delay time’ from an initial burst of star formation to the SN Ia explosion. For the SN Ia population as a whole, the distribution of these delay times – the delay time distribution (DTD) – is well described by a declining power law (Maoz & Gal-Yam, 2004; Totani et al., 2008; Maoz et al., 2010, 2014; Graur et al., 2015; Castrillo et al., 2021; Wiseman et al., 2021). When such a DTD is convolved with the stellar age distribution of stars across different galaxy types, a bimodal distribution for the age of SN progenitors at the time of explosion is observed (Mannucci et al., 2005; Scannapieco & Bildsten, 2005; Mannucci et al., 2006; Sullivan et al., 2006; Smith et al., 2012; Childress et al., 2014; Wiseman et al., 2021): so-called ‘prompt’ and ‘delayed ’ components.
@@
We fit the full dataset with this model by minimizing the negative log-likelihood, and show the data and best-fitting model in Fig 13. We verify that the two Gaussian model is preferred over a single Gaussian via the Akaike information criterion: taking the significance of the preference as $\exp{\Delta\mathrm{AIC}/2}$, we find that the double Gaussian is preferred to $6.8\sigma$. We do not use a likelihood ratio test since the models are not strictly nested. We find $\mu_{1,x_{1}}=0.22\pm 0.03$, $\sigma_{1,x_{1}}=0.67\pm 0.03$, $\mu_{2,x_{1}}=-0.86\pm 0.17$, $\sigma_{2,x_{1}}=0.80\pm 0.07$, and $a=0.37\pm 0.1$ which are broadly consistent with the results of Nicolas et al. (2021) and Ginolin et al. (2024a) except for $a$ which is significantly smaller than found in those works but consistent with the value for DES-SN5YR modelled in Wiseman et al. (2022). This small $a$ could be caused by our fixing of $y^{i}$ based on the galaxy colours: it is known that no galaxy observable perfectly traces the underlying properties, and even locally measured colours experience contamination whereby young SNe occur in populations measured to be old (Briday et al., 2022). If instead we fix the probability that any given SN is young to the average for our mean redshift according to the Rigault et al. (2020) drift model, $\delta(z=0.53)=0.67$, we find $\mu_{1,x_{1}}=0.06$, $\sigma_{1,x_{1}}=0.70$, $\mu_{2,x_{1}}=-1.65$, $\sigma_{2,x_{1}}=0.45$, and $a=0.58$. Note that these mean values are slightly more negative than the low-$z$ values from Nicolas et al. (2021) and Ginolin et al. (2024a) – this could be genuine redshift evolution or an effect of the use of SALT2 rather than SALT3 light curve fits in those works.
Next, we fit the inner and outer SNe separately, fixing the model parameters to those found using the aperture/annulus colour as a proxy for stellar population age, except for $a$ which we leave free. We find $a=0.32\pm 0.05$ in inner regions and $a=0.42\pm 0.04$ in outer regions, which means there may be a larger fraction of high-stretch mode SNe in red outer annuli compared to red central regions.
@@ 5 Summary
Using the DES five-year sample of photometrically-classified SNe Ia, we have analysed the relationships between projected and normalized host galaxy separation and the SN Ia light-curve properties and inferred cosmological distance measurements. Using the directional light radius (DLR) method as our normalisation of the galactocentric distances ($d_{\mathrm{DLR}}$), our main findings are:
  • We confirm previous findings of Toy et al. (2023) that the innermost regions of galaxies ($d_{\mathrm{DLR}}$$\leq 1$) host faster-declining SNe than the outer regions. We also confirm earlier results concerning SN Ia colour, and show that SNe Ia in the outer regions of galaxies ($d_{\mathrm{DLR}}>1.5$) are bluer ($\bar{c}=-0.023\pm 0.006$) than those at $d_{\mathrm{DLR}}\leq 1.5$ ($\bar{c}=0.010\pm 0.003$), a difference of $0.033\pm 0.007$ ($\simeq 4.7\sigma$).
  • We show, for the first time, that the difference in SN Ia post-standardization brightnesses between high and low-mass hosts reduces from $0.078\pm 0.011$ mag in the full sample to $0.036\pm 0.018$ mag for SNe Ia located in the outer regions of their host galaxies ($d_{\mathrm{DLR}}>1$), while increasing to $0.100\pm 0.014$ mag for SNe in the inner regions ($d_{\mathrm{DLR}}\leq 1$). The difference in the size of the mass step between inner and outer regions is $0.064\pm 0.023$ mag. The effect remains when splitting SNe Ia by their global galaxy $U-R$ colour, or by the $U-R$ colour in the inner aperture/outer annulus in which the SN occurred.
@@
  • There is no evidence that the outer regions of galaxies have dust laws affecting SNe Ia that change as a function of the galaxy’s stellar mass. Similarly, there is no evidence for an intrinsic luminosity difference between SNe Ia in the outer regions of low and high-mass galaxies.
@@
Calculating $d_{\mathrm{DLR}}$ is an algorithmically straight forward task, and selecting SNe Ia based on $d_{\mathrm{DLR}}$ reduces the need to account for host galaxy properties to standardize SN Ia brightnesses across different galaxies, without much computational effort. While we have not identified the underlying astrophysics that such projected galactocentric distances are tracing in the DES five-year SN Ia sample, we have shown that the standardized distance measurements from SNe Ia in the outer regions of galaxies have a reduced dependence on their global host galaxy properties.
Restricting a cosmological analysis to SNe Ia in the outer regions of their host galaxies reduces the sample size to around a third. This will increase the statistical uncertainties. However, the reduction in the astrophysical systematic uncertainties (and complications) gained from using such a sample, coupled with the very large sample sizes expected in future experiments such as the Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), means that such a selection is likely to be a useful validation of the results found using a full sample.
@@ Acknowledgements
We thank the referee for a thoughtful review. All authors have contributed to the drafting of this manuscript. MT devised the project and led the analysis. PW and MS provided scientific guidance and support with the production of the manuscript. MS contributed host galaxy fitting and wrote substantial sections of the manuscript, and PW ran 1D bias corrections and provided the colour and stretch analyses. DS and MV internally reviewed the work and provided extensive feedback. CF, CL, JL, LK, LG, RK, PS, and TD provided comments on the analysis and interpretation. All aforementioned authors as well as AM, BP, BS, DB, and MS contributed to the DES-SN5YR data and methods used in this paper. The remaining authors have made contributions to this paper that include, but are not limited to, the construction of DECam and other aspects of collecting the data; data processing and calibration; developing broadly used methods, codes, and simulations; running the pipelines and validation tests; and promoting the science analysis.
@@
Appendix A BBC1D distances
In Table 5 we present the ‘BBC1D’ corrected distance moduli for our 1531 DES5YR SNe. A full version of the table can be found online. Full data for each SN can be found in the DES5YR data release described in Sánchez et al. (2024) and found at https://des-sn-dr.readthedocs.io/en/latest/.
CIDa $\mu^{\mathrm{b}}$ $\Delta\mu^{\mathrm{c}}$ $\mu_{\mathrm{bias}}^{\mathrm{d}}$
1702083 41.7034 0.1041 -0.0050
1435094 43.6092 -0.5358 -0.0697
1309227 44.2862 0.3922 -0.0472
1324542 42.5540 0.1201 -0.0325
1320474 43.6429 0.2471 -0.1081
Table 5: a) DES unique SN identifier; b) Distance modulus from Eq. 1; c) Residual of $\mu$ with respect to the fiducial $\Lambda$CDM cosmological model (Eq 2); d) Size of the 1D bias correction applied to achieve the $\mu$ estimate.
1 School of Physics and Astronomy, University of Southampton, Southampton, SO17 1BJ, UK
2 Department of Physics, Duke University Durham, NC 27708, USA
3 Einstein Fellow
4 Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH, United Kingdom
5 Center for Astrophysics |||| Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
6 School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia
7 Institut d’Estudis Espacials de Catalunya (IEEC), 08034 Barcelona, Spain
8 Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, 08193 Barcelona, Spain
9 Centre for Gravitational Astrophysics, College of Science, The Australian National University, ACT 2601, Australia
10 The Research School of Astronomy and Astrophysics, Australian National University, ACT 2601, Australia
11 Department of Physics and Astronomy, University of Pennsylvania, Philadelphia, PA 19104, USA
12 Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK
13 Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637, USA
14 Kavli Institute for Cosmological Physics, University of Chicago, Chicago, IL 60637, USA
15 Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Victoria 3122, Australia
16 Université Grenoble Alpes, CNRS, LPSC-IN2P3, 38000 Grenoble, France
17 Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
18 Physics Department, Lancaster University, Lancaster, LA1 4YB, UK
19 Laboratório Interinstitucional de e-Astronomia - LIneA, Av. Pastor Martin Luther King Jr, 126 Del Castilho, Nova América Offices, Torre 3000/sala 817 CEP: 20765-000, Brazil
20 Fermi National Accelerator Laboratory, P. O. Box 500, Batavia, IL 60510, USA
21 Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
22 Kavli Institute for Particle Astrophysics & Cosmology, P. O. Box 2450, Stanford University, Stanford, CA 94305, USA
23 SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
24 Instituto de Astrofisica de Canarias, E-38205 La Laguna, Tenerife, Spain
25 Universidad de La Laguna, Dpto. Astrofísica, E-38206 La Laguna, Tenerife, Spain
26 Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain
27 Hamburger Sternwarte, Universität Hamburg, Gojenbergsweg 112, 21029 Hamburg, Germany
28 Department of Physics, IIT Hyderabad, Kandi, Telangana 502285, India
29 California Institute of Technology, 1200 East California Blvd, MC 249-17, Pasadena, CA 91125, USA
30 Institute of Theoretical Astrophysics, University of Oslo. P.O. Box 1029 Blindern, NO-0315 Oslo, Norway
31 Instituto de Fisica Teorica UAM/CSIC, Universidad Autonoma de Madrid, 28049 Madrid, Spain
32 Center for Astrophysical Surveys, National Center for Supercomputing Applications, 1205 West Clark St., Urbana, IL 61801, USA
33 Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green Street, Urbana, IL 61801, USA
34 Santa Cruz Institute for Particle Physics, Santa Cruz, CA 95064, USA
35 Center for Cosmology and Astro-Particle Physics, The Ohio State University, Columbus, OH 43210, USA
36 Department of Physics, The Ohio State University, Columbus, OH 43210, USA
37 Australian Astronomical Optics, Macquarie University, North Ryde, NSW 2113, Australia
38 Lowell Observatory, 1400 Mars Hill Rd, Flagstaff, AZ 86001, USA
39 Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA
40 George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, and Department of Physics and Astronomy, Texas A&M University, College Station, TX 77843, USA
41 LPSC Grenoble - 53, Avenue des Martyrs 38026 Grenoble, France
42 Institució Catalana de Recerca i Estudis Avançats, E-08010 Barcelona, Spain
43 Department of Physics, Carnegie Mellon University, Pittsburgh, Pennsylvania 15312, USA
44 Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil
45 Department of Physics and Astronomy, Pevensey Building, University of Sussex, Brighton, BN1 9QH, UK
46 Department of Physics, Northeastern University, Boston, MA 02115, USA
47 Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Madrid, Spain
48 Physik-Institut, University of Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland
49 Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831
50 High Energy Physics Division, Argonne National Laboratory, 9700 S. Cass Avenue, Argonne, IL 60439, USA
51 Cerro Tololo Inter-American Observatory, NSF’s National Optical-Infrared Astronomy Research Laboratory, Casilla 603, La Serena, Chile
52 Department of Astronomy, University of California, Berkeley, 501 Campbell Hall, Berkeley, CA 94720, USA
53 Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA